OpenStudy (princessaurora):

Write a finite geometric series representing the total amount of water leaked after h hours. *I can't post a pic of the beaker but the fill line is at 75ml and the beaker goes 10,20,30,40,50,60,70,80,90,100*

OpenStudy (campbell_st):

so you you know any other information...?

OpenStudy (princessaurora):

The top of the page says, " Neil has a leaky faucet in his bathroom. He decides to find the amount of water leaking from the faucet over a period of time. For the first hour, he measured the volume of water that leaked by collecting it in a beaker."

OpenStudy (campbell_st):

ok... so what is the volume of water for the hour...?

OpenStudy (princessaurora):

The volume of water collected in the beaker in the first hour would be 75ml.

OpenStudy (campbell_st):

ok... so the leak is 75ml/h so what other information do you have..?

OpenStudy (princessaurora):

For the second hour the volume is 81.25 ml.

OpenStudy (campbell_st):

what about the 3rd hour...?

OpenStudy (princessaurora):

It doesn't tell me but it says the amount of water leaking from the faucet every hour increases uniformly by (1/12) of the amount leaked in the previous hour.

OpenStudy (campbell_st):

ok... so you have the initial term 75 common ratio 13/12 so does 81.15 = 75 x 13/12..?

OpenStudy (princessaurora):

Is that a geometric series though?

OpenStudy (campbell_st):

It will be  so the initial volume is 75 ml the common ratio is 13/12 so a turn in the series is \[T_{n} = 75 \times (\frac{13}{12})^{n - 1}\]